Results from MathSciNet: Mathematical Reviews on the Web \begin{matrix}c \\ ◯ \end{matrix} Copyright American Mathematical Society 1999

99f:20001a 20-02 20Cxx 20J06

Benson, D. J. [Benson, David John] ^{(}1-GA^{)}

⋆Representations and cohomology. I. (English. English summary) Basic representation theory of finite groups and associative algebras. Second edition. Cambridge Studies in Advanced Mathematics, 30. Cambridge University Press, Cambridge, 1998. xii+246 pp. $29.95. ISBN 0-521-63653-1

99f:20001b 20-02 19-02 20Cxx 20Jxx 55-02

Benson, D. J. [Benson, David John] ^{(}1-GA^{)}

⋆Representations and cohomology. II. (English. English summary) Cohomology of groups and modules. Second edition. Cambridge Studies in Advanced Mathematics, 31. Cambridge University Press, Cambridge, 1998. xii+279 pp. $29.95. ISBN 0-521-63652-3 From the introduction: “In preparing the paperback edition, I have taken the liberty of completely retypesetting the book using the enhanced features of L^{A}TEX 2_{ε}, AMS-L^{A}TEX 1.2 and XY-pic 3.5. Apart from this, I have corrected those errors of which I am aware.

“The most extensively changed sections are Sections 2.2 and 3.1 of Volume I and Section 5.8 of Volume II, which contained major flaws in the original edition. In addition, in Section 3.1 of Volume I, I have changed to the more usual definition of Hopf algebra in which an antipode is part of the definition, reserving the term bialgebra for the version without an antipode.

“I have made every effort to preserve the numbering of the sections, theorems, references, and so on from the first edition, in order to avoid reference problems. The only exception is that in Volume I, Definition 3.1.5 has disappeared and there is now a Proposition 3.1.5. I have also updated the bibliography and improved the index”

{The first edition has been reviewed [Vol. I; MR 92m:20005; Vol. II; MR 93g:20099].}